Hastings Corporation, a large telecommunications company, is considering the purchase of Vandell Corporation, a smaller, regional provider in the same industry. Hastings’ management foresees much synergy between the two firms as a key outcome of a possible merger. As the Acquisitions Analyst employed by Hastings, you have developed the following estimates for Vandell’s incremental “Free Cash Flows” (FCF) for the upcoming four years: Year Free Cash Flow 1 $2.5 million 2 $2.9 million 3 $3.4 million 4 $3.57 million These “Free Cash Flow” values for Vandell Corporation include all acquisition effects. Your analysis further predicts that “Free Cash Flows” occurring after Year 4 will grow at a constant four percent (4%) annual rate. Vandell Corporation is estimated to have a “Weighted Average Cost of Capital” (WACC) of 8%, based on the capital structure that will be put in place following the acquisition, if it is made. During your employment at Hastings Corporation, you have learned that the management team, led by Don Hastings, prefers the “Corporate Valuation Model” (CVM) method when evaluating possible acquisition candidates. You are preparing a presentation on this proposal for the management meeting tomorrow; the possible acquisition of Vandell Corporation is considered to be the most important agenda item for this meeting. In preparation for this meeting, you prepare answers to the following questions: Calculate the “Horizon Value” (HV) as of the end of Year Four for Vandell Corporation’s expected “Free Cash Flows” that will occur after that date. What is the present value of the expected “Free Cash Flows” that Vandell Corporation is projected to generate, including the “Horizon Value,” given the assumptions outlined above? If Hastings Corporation acquires only the assets and operations of Vandell Corporation, what is the maximum offer that Hastings should make when negotiating the purchase of Vandell? a. Why should Hastings not pay more than this amount for Vandell? Now assume that Vandell Corporation has $30 million of debt outstanding. If Hastings Corporation were to acquire the stock of Vandell Corporation (i.e., an equity acquisition), what would then be the maximum amount that Hastings should pay for Vandell (i.e., what is Vandell Corporation’s “Net Asset Value?”)?
Answer #1
[$ in millions] 0 1 2 3 4
FCF $                 2.50 $              2.90 $                 3.40 $                3.57
PVIF at 8% [PVIF = 1/1.08^t) 1 0.92593 0.85734 0.79383 0.73503
PV at 8% $                 2.31 $              2.49 $                 2.70 $                2.62
Sum of PVs of FCF--t1 to t4 $              10.12
Horizon value of FCF = 3.57*1.04/(0.08-0.04) = $              92.82
PV of horizon value = 92.82/1.08^4 = $              68.23
Total value of FCF $              78.35
ANSWERS:
1] HV at the end of year four $              92.82
2] PV of expected FCF [including HV] $              78.35
3] Max. that Hastings should pay for assets and operation Vandell = $              78.35
4] Max. that Hastings should pay for equity [NAV] of Vandell = 78.35-30 = $              48.35